What is the speed of a satellite moving in a circular orbit at a height of 3800km above

the surface of the earth? (b) what is the period of the satellite in hours? The mass of the

earth is 5.97x10^24 kg. The radius of the earth is 6.38x10^6 m.


1
Expert's answer
2021-10-21T16:36:05-0400

Given:

"G=6.67*10^{-11}\:\rm N\cdot m^2/kg^2"

"R_E=6.38*10^{6}\:\rm m"

"h=3.80*10^{6}\:\rm m"

"M_E=5.97*10^{24}\:\rm kg"


(a) the speed of a satellite moving in a circular orbit 

"v=\sqrt{\frac{GM_E}{R_E+h}}"

"v=\sqrt{\frac{6.67*10^{-11}*5.97*10^{24}}{6.38*10^{6}+3.80*10^{6}}}=6250\:\rm m/s"

(b) what is the period of the satellite in hours

"T=\frac{2\pi(R_E+h)}{v}"

"T=\frac{2\pi(6.38*10^{6}+3.80*10^{6})}{6250}=10200\:\rm s=2.84\: hr"